OpenStudy (anonymous):

HELP ME PLEASE , I WILL MEDAL Solve, using the substitution method. y – 2x = 8 16 + 4x = 2y A. There are an infinite number of solutions. B. The only solution is (1, 10). C. There is no solution. D. The only solution is (24, 0).

OpenStudy (anonymous):

@barreraA

OpenStudy (barreraa):

C. There is no solution. @jayonna15

OpenStudy (anonymous):

thank you

OpenStudy (barreraa):

y - 2x = 8 --- we want to get y by itself, therefore, we want the -2x on the other side. To do this, we do the opposite. It is showing -2x, so we add 2x to both sides. y - 2x + 2x = 2x + 8 -- combine like terms y = 2x + 8 now that we have y = 2x + 8, we can sub 2x + 8 in for y in the other equation. 16 + 4x = 2y 16 + 4x = 2(2x + 8) -- distribute through the parenthesis 16 + 4x = 4x + 16 -- subtract 4x from both sides 16 + 4x - 4x = 4x - 4x + 16 -- combine like terms 16 = 16 if the solution ends like this, there are INFINITE SOLUTIONS if it would have ended like this : 16 = 14...then it would be no solutions.

OpenStudy (barreraa):

and np @jayonna15 anytime